# Inference on many jumps in nonparametric panel regression models

### v2

@@ Abstract
We investigate the significance of change-point or jump effects within fully nonparametric regression contexts, with a particular focus on panel data scenarios where data generation processes vary across individual or group units, and error terms may display complex dependency structures. In our setting the threshold effect depends on a specific covariate, and we permit the true nonparametric regression to vary based on additional latent variables. We propose two uniform testing procedures: one to assess the existence of change-point effects and another to evaluate the uniformity of such effects across units. Even though the underlying data generation processes are neither independent nor identically distributed, our approach involves deriving a straightforward analytical expression to approximate the variance-covariance structure of change-point effects under general dependency conditions. Notably, when Gaussian approximations are made to these test statistics, the intricate dependency structures within the data can be safely disregarded owing to the localized nature of the statistics. This finding bears significant implications for obtaining critical values. Through extensive simulations, we demonstrate that our tests exhibit excellent control over size and reasonable power performance in finite samples, irrespective of strong cross-sectional and weak serial dependency within the data. Furthermore, applying our tests to two datasets reveals the existence of significant nonsmooth effects in both cases.
Keywords: high-dimensional time series; temporal and cross-sectional dependence; changepoint analysis; Gaussian approximation; simultaneous tests.
@@ 1 Introduction
Recently, there has been a notable surge in the study and application of changepoint analysis in high-dimensional time series data (see, e.g., Cho and Fryzlewicz, (2015), Cao et al., (2018), Chen et al., (2022), and Li et al., (2024), along with the references cited therein). The focus of changepoint analysis has shifted from the traditional study of temporal changepoints to exploring changepoints within nonlinear regression models. Detecting changepoints in regression models is akin to clustering data points based on similar characteristics or behaviors. Chang et al., (2015) reviews several intriguing applications of changepoint regression. In the medical domain, changepoint analyses often indicate shifts in the impact of certain risk factors on outcomes (see, e.g., Pastor and Guallar, (1998) and Muggeo, (2003)). Additionally, there is significant interest in detecting and evaluating changepoint regressions within econometrics. For example, Lee, (2008) studies the effect of party incumbency in U.S. House Elections. The purpose of these studies is to estimate not only where the threshold effect exists but also the magnitude of the effects.
@@
There is extensive literature on parametric high-dimensional changepoint regression analysis. Wang et al., (2021), Cho et al., (2024), and Rinaldo et al., (2021) investigate changepoint analysis of coefficients in high-dimensional linear regression models using penalization methods. Lee et al., (2016) examines a high-dimensional regression model with a changepoint due to a covariate threshold in a cross-sectional study. These studies primarily focus on parametric models with high-dimensional covariates. Similar to Lee et al., (2016), this paper also considers changepoint regression due to a covariate threshold. However, our focus is on nonparametric regression with a high-dimensional changepoint effect driven by heterogeneity in the panel data structure. Specifically, we investigate the heterogeneous nonparametric panel data with jumps in the conditional mean functions:
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt},$ (1.1)
$\mathbb{E}(e_{jt}|X_{jt})=0,$ (1.2)
where $Y_{jt}$ and $X_{jt}$ are the outcome and covariate variables for entity $j\in[N]:=\{1,\ldots,N\}$ at time point $t\in[T],$ $f_{j}(\cdot)$ is a continuous function, $c_{0j}$ is a threshold value that is either known or unknown (if $\gamma_{j}=0$, then without loss of generality set $c_{0j}=0$ for the unknown case), $\gamma_{j}$ is the jump effect at $c_{0j}$ and $e_{jt}$ is the error term. When $c_{0j}$’s are known, the above model is linked to a panel version of the regression discontinuity model that is widely studied in an i.i.d. setup. Our primary focus is to test whether changepoints exist and if the jump effects are consistent across different entities under general assumptions about $e_{jt}$. Specifically, we consider $e_{jt}$ as a reduced-form error term with a complex structure that accounts for latent variables $U_{jt}$, which can be arbitrarily correlated with $X_{jt}$. The detailed settings for $e_{jt}$ are discussed in Section 2.1.
@@
Our paper is related to four strands of literature in statistics and econometrics. First, we add to the extensive body of work on changepoint analysis and threshold regression models. Given the vast literature in this area, we only review briefly cases involving high-dimensional data. For example, Li et al., (2016) consider estimation of panel data models with interactive fixed effects (IFEs) and multiple structure changes, and Manner et al., (2019) consider testing for structural breaks in factor copula models. Miao et al., 2020a and Miao et al., 2020b consider the estimation of a panel threshold regression model with IFEs and latent group structures, respectively; Massacci and Trapani, (2022) study inference for high-dimensional threshold regressions with common stochastic trends; Barassi et al., (2023) study a heterogeneous panel threshold regression model with IFEs. Second, our paper is related to the literature on testing homogeneity in a panel data setting. For example, Phillips and Sul, (2003), Pesaran and Yamagata, (2008) and Su and Chen, (2013) propose tests for slope coefficient heterogeneity in a linear regression model with or without cross-sectional dependence. In the presence of slope heterogeneity, Su et al., (2016) and Su and Ju, (2018) model the slope heterogeneity via latent group structures. Third, in terms of methodology, our paper aligns with research on nonparametric trends in multivariate or high-dimensional time series. Inference on smooth trend functions in nonparametric models for time series data is a well-established field. For example, Zhang and Wu, (2012) study inference of time-varying regression models, Karmakar et al., (2022) study simultaneous inference for time-varying models, and Gao et al., (2024) consider estimation and inference for time-varying multivariate causal processes. Extending the above works to high-dimensional settings, Chen and Wu, (2018) consider statistical inference for trends in high-dimensional time series. Finally, our paper is connected to the extensive literature on regression discontinuity design (RDD). Our approach can be extended to a panel RDD framework with specific design assumptions, particularly addressing individual effects that may cause biased RDD estimators or result in efficiency loss when estimating homogeneous treatment effects in finite samples. Different from RDD, our work focuses on inference of jumps in the nonparametric regression functions for high-dimensional data models and is thus also related to a set of literature regarding inference of jumps in nonparametric trends. However, the existing literature concentrates only on the cross-sectional case without dependency; see, e.g., Qiu, (1998), Müller and Stadtmüller, (1999) and Spokoiny, (1998) who focus on the estimation of discontinuous regression functions in the one-dimensional case based on cross-sectional data. Another distinct feature is that we explicitly handle the case with missing covariates or models with misspecified variables. Theoretically, validating such an inference procedure requires working with a reduced-form model that has specific variance-covariance structures. This aspect of theoretical validity has not been explored in the time-varying trend literature. Establishing the validity of uniform inference is a major theoretical contribution that sets our work apart in the study of varying trends.
The methodological contributions are twofold. First, we develop algorithms to test the overall significance or heterogeneity of threshold effects and examine the asymptotic properties of the resulting test statistics. When $N$ is large, assuming homogeneous threshold effects $\gamma_{j}$ across all $j$ is restrictive, and inferences based on this assumption can be misleading if the threshold effects are heterogeneous. Therefore, it is important to test for heterogeneous threshold effects. Given the well-documented cross-sectional and serial dependence in panel data, it is crucial to develop tests that account for both types of dependence. Specifically, we develop a supremum-type test that can distinguish between a constant or zero average changepoint effect and the alternative that different individuals exist for different individuals/groups. The test is built upon the supremum of $N$ kernel-estimator-based test statistics, each of which is constructed from a time-series local linear regression. Due to the use of local smoothing, neither the weak dependence along the time dimension nor the weak or strong dependence along the individual/group dimension enters the asymptotic distribution of the final statistic. Interestingly, we find strong cross-sectional dependence is allowed in our error terms. As a practical recommendation, we find that under fairly weak assumptions, if there are clustering dependency patterns in $e_{jt}$, the correlation in the estimated statistics can be safely ignored, and only adjustments for heteroskedasticity are needed.
Second, we extend existing Gaussian approximation (GA) results to accommodate latent variables in the error terms. The theoretical challenges arise when the number of groups $N$ is much larger than $Tb_{j}$, the effective sample size, where $b_{j}$ represents a certain bandwidth parameter. This high dimensionality makes it difficult to apply the standard central limit theorem. Our theoretical results extend the existing GA results from Chernozhukov et al., (2015), Chernozhukov et al., 2017a , Zhang and Wu, (2017) and Chernozhukov et al., (2019). Our test setup is similar to that in Chernozhukov et al., (2019), but the main difference is that our test statistic is nonparametric, involving the differences between two kernel estimators. The primary technical challenge is accounting for underlying latent variables in our data generating process (DGP). To address this, we introduce conditioning arguments to derive new high-dimensional GA results that accommodate general spatial and temporal dependency structures in the error terms. These GA results are crucial for determining the critical values of our supremum test.
On the empirical side, we employ our methods in two empirical applications. First, we detect jumps in the news impact curve for constituents in the S&P 500 stock index with an unknown threshold. In the second application, we study the effect of party incumbency as determined by the 50% threshold in the two-party vote share in the US House Elections. Our findings confirm the presence of a significant positive incumbency effect as studied by Lee, (2008), and we detect significant heterogeneity in the threshold effect across the US states.
@@
The rest of the paper is organized as follows. Section 2 provides the model setup and the practical steps of the testing procedure. Section 3 presents the assumptions and delivers the main theorems. In Section 4 we provide some Monte Carlo simulation results. Section 5 concludes. The Online Supplement contains the proofs of all results, some additional simulation results, and empirical applications.
@@
Notation. For a vector $v=(v_{1},\ldots,v_{d})\in\mathbb{R}^{d}$ and $q>0$, we denote $|v|_{q}=(\sum_{i=1}^{d}|v_{i}|^{q})^{1/q}$ and $|v|_{\infty}=\max_{1\leq i\leq d}|v_{i}|$. For a matrix $A=(a_{i,j})_{1\leq i\leq m,1\leq j\leq n}$, we define the max norm $|A|_{\text{max}}=\max_{i,j}|a_{i,j}|$. For $s>0$ and a random vector $X$, we say $X\in\mathcal{L}^{s}$ if $\lVert X\rVert_{s}=[\mathbb{E}(|X|^{s})]^{1/s}<\infty$. For two positive number sequences $(a_{n})$ and $(b_{n})$, we say $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ (resp. $a_{n}\asymp b_{n}$) if there exists $C>0$ such that $a_{n}/b_{n}\leq C$ (resp. $1/C\leq a_{n}/b_{n}\leq C$) for all large $n$, and say $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\rightarrow 0$ as $n\rightarrow\infty$. We set $(X_{n})$ and $(Y_{n})$ to be two sequences of random variables. Write $X_{n}=O_{\mathbb{P}}(Y_{n})$ if for any $\epsilon>0$, there exists $C>0$ such that $\mathbb{P}(X_{n}/Y_{n}\leq C)>1-\epsilon$ for all large $n$, and say $X_{n}=o_{\mathbb{P}}(Y_{n})$ if $X_{n}/Y_{n}\rightarrow 0$ in probability as $n\rightarrow\infty$. We denote the centered random variable $X$ by $\mathbb{E}_{0}(X)$, i.e., $\mathbb{E}_{0}(X)=X-\mathbb{E}(X)$.
@@ 2.1 Model Setup
In this subsection, we formulate our model. The data $(Y_{jt},X_{jt})_{j\in[N],\ t\in[T]}$ are collected for individual/group $j$ at time $t$, where $Y_{jt}$ is the outcome variable and $X_{jt}$ is a continuous covariate variable. We work with the following full model:
where $U_{jt}$ is a latent random variable that can be arbitrarily correlated with $X_{jt},$ $\epsilon_{jt}$ is a mean zero term that is independent of $\left(X_{jt},U_{jt}\right)$. Further, $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ are continuous functions of $X_{jt}$ and $U_{jt}$ for the baseline effect, the jump effect and the volatility, respectively. The motivation for considering such a model with a latent variable $U_{jt}$ is that, in practice, the outcome is often affected by unobserved variables and factors. For example, in our application (Section E.1 in the online supplement), we analyze the threshold effect of asset returns. $Y_{jt}$ is the stock volatility, $X_{jt}$ is the lagged stock return, and $U_{jt}$ are unobserved risk factors. Modeling $U_{jt}$ is important in this circumstance as the volatility of a single stock is most likely affected by other risk factors such as news events or unobserved market factors.
This paper aims to explore methods for conducting simultaneous inference on threshold effect estimators in a large $N$ and large $T$ setup. In particular, we allow $N\gg T.$ Our primary focus is on the threshold effect, starting with the case where the threshold value $c_{0j}$ is known and then extending the analysis to the unknown case in Subsections 2.3 and 3.4.
Assuming that $X_{jt}$ is a continuous random variable, the jump effect can be identified as a “gap” in the conditional expectations of the dependent variable at the threshold:
$\gamma_{j}=\lim_{x_{j}\downarrow c_{0j}}\mathbb{E}(Y_{jt}|X_{jt}=x_{j})-\lim_{x_{j}\uparrow c_{0j}}\mathbb{E}(Y_{jt}|X_{jt}=x_{j}).$ (2.2)
Let $\tilde{\tau}_{j}(X_{jt})=\mathbb{E}[\tau_{j}(X_{jt},U_{jt})|X_{jt}]$ and $\tilde{h}_{j}(X_{jt})=\mathbb{E[}h_{j}(X_{jt},U_{jt})|X_{jt}]$. Then the conditional expectation of the observed outcome given $X_{jt}$ is $\mathbb{E}(Y_{jt}|X_{jt})=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}.$  Note that although model (2.1) takes a complicated structure it can be rewritten into model (1.1) after some simple transformation. Specifically denote $\gamma_{j}=\tilde{\tau}_{j}(c_{0j}),$
$f_{j}(X_{jt})=\tilde{h}_{j}(X_{jt})+[\tilde{\tau}_{j}(X_{jt})-\tilde{\tau}_{j}(c_{0j})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$
$\varepsilon_{j}(X_{jt},U_{jt})=h_{j}(X_{jt},U_{jt})-\tilde{h}_{j}(X_{jt})+[\tau_{j}(X_{jt},U_{jt})-\tilde{\tau}_{j}(X_{jt})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}\text{ and}$
@@
Therefore model (2.1) is our full model and can be reduced to the working model in Equation (1.1). The only technical difficulty this introduces to our framework is the need to handle the special variance structure of the error term. It is easy to verify that $\mathbb{E}[\varepsilon_{j}(X_{jt},U_{jt})|X_{jt}]=0$ which further leads to $\mathbb{E}[e_{jt}|X_{jt}]=0.$ Note that $f_{j}\left(\cdot\right)$ is continuous provided $\tilde{h}_{j}\left(\cdot\right)$ and $\tilde{\tau}_{j}(\cdot)$ are. So $f_{j}(X_{jt})$ and $\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ correspond to the continuous and discrete part of $\mathbb{E}(Y_{jt}|X_{jt}),$ respectively. The noise term $e_{jt}$ has a complex structure, influenced both by an unobserved random variable and by the heterogeneous volatility of the observations. This intricate structure, combined with the temporal and cross-sectional dependencies of the noise, makes the inference process particularly challenging.
Remark 1.
$Y_{jt}=\alpha_{j}+\nu_{t}+f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$ (2.4)
$Y_{jt}=\tilde{f}_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\tilde{e}_{jt}.$ (2.5)
Remark 2.
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}+Z_{jt}^{\top}\theta.$ (2.6)

### v3

@@ Abstract
We investigate the significance of change-points within fully nonparametric regression contexts, with a particular focus on panel data where data generation processes vary across units, and error terms may display complex dependency structures. In our setting the threshold effect depends on one specific covariate, and we permit the true nonparametric regression to vary based on additional (latent) variables. We propose two uniform testing procedures: one to assess the existence of change-points and another to evaluate the uniformity of such effects across units. Our approach involves deriving a straightforward analytical expression to approximate the variance-covariance structure of change-point effects under general dependency conditions. Notably, when Gaussian approximations are made to these test statistics, the intricate dependency structures within the data can be safely disregarded owing to the localized nature of the statistics. This finding bears significant implications for obtaining critical values. Through extensive simulations, we demonstrate that our tests exhibit excellent control over size and reasonable power performance in finite samples, irrespective of strong cross-sectional and weak serial dependency within the data. Furthermore, applying our tests to two datasets reveals the existence of significant nonsmooth effects in both cases.
Keywords: high-dimensional time series; temporal and cross-sectional dependence; changepoint analysis; Gaussian approximation; simultaneous tests; threshold regression.
@@ 1 Introduction
Recently, there has been a notable surge in the study and application of changepoint analysis in high-dimensional time series data (see, e.g., Cho and Fryzlewicz, (2015), Cao et al., (2018), Chen et al., (2022), and Li et al., (2024), along with the references cited therein). Detecting changepoints in regression models is important and has wide applications in the field of biology (Pastor and Guallar, , 1998), agricultural sciences (Freeman et al., , 1998), econometrics (Li et al., , 2024), etc. The purpose of these studies is to estimate not only where the threshold effect exists but also the magnitude of the effects.
@@
There is extensive literature on parametric high-dimensional changepoint regression analysis. Li et al., (2016), Manner et al., (2019) Wang et al., (2021), Rinaldo et al., (2021), and Cho et al., (2024) investigate changepoint analysis of coefficients in high-dimensional linear regression models using penalization methods. Lee et al., (2016) examine a high-dimensional regression model with a changepoint due to a covariate threshold in a cross-sectional study. These studies majorly focus on parametric models with high-dimensional covariates. Similar to Lee et al., (2016), this paper also considers changepoint regression due to a covariate threshold. However, our focus is on nonparametric regression with a high-dimensional changepoint effect driven by heterogeneity in the panel data structure. Specifically, we investigate the heterogeneous nonparametric panel data with jumps in the conditional mean functions:
where $X_{jt}$ is a one-dimensional threshold variable, $Y_{jt}\in\mathbb{R}^{1}$ and $U_{jt}\in\mathbb{R}^{d}$ are the outcome and covariate variables that can be observed or unobserved for individual $j\in[N]:=\{1,\ldots,N\}$ at time point $t\in[T],$ $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ are continuous functions, $c_{0j}$ is a threshold value that is either known or unknown and $\epsilon_{jt}$ is the error term. Our main focus is to test whether changepoint effects related to $\tau_{j}(X_{jt},U_{jt})$ exist and to determine if the jump sizes are identical uniformly across individuals $j$ under general spatial temporal dependence assumptions.
As a motivating example for our model, we explore an application involving stock lagged returns and volatility using threshold regression. Threshold regression, dating back to Chan et al., (1985) and more recently studied in works such as Duffy et al., (2023), is mainly formulated within a linear parametric setting. Our approach naturally extends this framework to the non-parametric setting. Specifically, in our first application, revising the news impact curve of Engle and Ng, (1993), $Y_{jt}$ is the volatility of firm $j$’s stock at time $t$ and $X_{jt}$ is the lagged stock return. A trader would like to discover whether there are asymmetric effects of $X_{jt}$ on $Y_{jt}$, as well as whether and where a discontinuous relationship occurs, for trading and risk management purposes. If there are multiple stocks jump, are they all identical uniformly? In Section 5 in our empirical section, we identify significant threshold effects, demonstrating that our methodology is crucial for predicting stock volatility and effectively managing risk. In another application, we study the effect of party incumbency as determined by the 50% threshold in the two-party vote share in the US House Elections. Our findings confirm the presence of a significant positive incumbency effect as studied by Lee, (2008), and we detect significant heterogeneity in the threshold effect across the US states.
The methodological contributions are twofold. First, we develop algorithms to test for overall threshold effects, examine the potential heterogeneity effects and derive the asymptotic properties of the resulting test statistics. Due to the use of local smoothing, neither the weak dependence along the time dimension nor the weak or strong dependence along the individual/group dimension enters the asymptotic distribution of the final statistic. Interestingly, we find strong cross-sectional dependence is allowed in our error terms. As a practical recommendation, we find that under fairly weak assumptions, if there are clustering dependency patterns in $e_{jt}$, the correlation in the estimated statistics can be safely ignored, and only adjustments for heteroskedasticity are needed. Second, the standard central limit theorem is challenging to apply to high-dimensional time series when $N$ significantly exceeds the effective sample size. Previously, Chernozhukov et al., (2015); Chernozhukov et al., 2017a , Zhang and Wu, (2017), Chernozhukov et al., (2019), and others addressed this issue by introducing high-dimensional Gaussian approximation (GA) methods. However, the existing literature cannot address our problem. The key difficulty is to incorporate possibly non-identically distributed time series under general spatial and temporal dependency structures and more importantly latent variables. In this paper, we propose a new high dimensional GA for weighted statistics that address all these issues. These GA results are essential for determining the critical values of our supremum test.
@@
Our paper is related to three strands of literature in statistics. First, we add to the extensive body of work on threshold regression for high dimensional data. For example, Miao et al., 2020a and Miao et al., 2020b consider the estimation of a panel threshold regression model with interactive fixed effects (IFEs) and latent group structures, respectively; Massacci and Trapani, (2022) study inference for high-dimensional threshold regressions with common stochastic trends. The aforementioned literature does not account for potential heterogeneous threshold effects. Ignoring these effects may bias estimators or lead to efficiency loss when estimating homogeneous threshold effects in finite samples. Second, our paper contributes to the literature on testing homogeneity in a panel data setting. For example, Phillips and Sul, (2003), Pesaran and Yamagata, (2008) and Su and Chen, (2013) propose tests for slope coefficient heterogeneity in a linear regression model with or without cross-sectional dependence. In the presence of slope heterogeneity, Su et al., (2016) and Su and Ju, (2018) model the slope heterogeneity via latent group structures. Barassi et al., (2023) study a heterogeneous panel threshold regression model with IFEs. Above literature all focus on parametric setting. Our work complements these research streams by adopting nonparametric approaches, which address size and power distortions caused by model misspecifications. Third, our nonparametric approach is related with literature in trends estimation. Inference on smooth trend functions in nonparametric models for time series data is a well-established field; see Zhang and Wu, (2012), Karmakar et al., (2022), Gao et al., (2024), Chen and Wu, (2018), among others. Distinctly, while the above literature examines nonlinear dynamic models, they does not address discontinuity in trends, particularly heterogeneous jumps. The nonparametric literature on jumps in trends mostly focuses on independent and identically distributed (i.i.d.) settings, without accounting for heterogeneity or high dimensionality; see, for example, Qiu, (1998), Müller and Stadtmüller, (1999), and Spokoiny, (1998).
Another distinct feature is that we explicitly handle the case with missing covariates or models with misspecified variables. Theoretically, validating such an inference procedure requires working with a reduced-form model that has specific variance-covariance structures. This aspect of theoretical validity has not been explored in the time-varying trend literature. Establishing the validity of uniform inference is a major theoretical contribution that sets our work apart in the study of varying trends.
@@
The rest of the paper is organized as follows. Section 2 provides the model setup and the practical steps of the testing procedure. Section 3 presents the assumptions and delivers the main theorems. In Section 4 we provide simulation results. [^1] The code for our method is available in the R package hdthreshold (https://cran.r-project.org/web/packages/hdthreshold/index.html), with an illustration accessible at https://rpubs.com/Lk1110/1259698. Real data applications are discussed in Section 5, and the paper concludes in Section 6. The Supplementary Material includes additional theorems, proofs of all results, extra simulation outcomes, and tables for the empirical applications.
@@
Notation. For a vector $v=(v_{1},\ldots,v_{d})\in\mathbb{R}^{d}$ and $q>0$, we denote $|v|_{q}=(\sum_{i=1}^{d}|v_{i}|^{q})^{1/q}$ and $|v|_{\infty}=\max_{1\leq i\leq d}|v_{i}|$. For a matrix $A=(a_{i,j})_{1\leq i\leq m,1\leq j\leq n}$, we define the max norm $|A|_{\text{max}}=\max_{i,j}|a_{i,j}|$. For $s>0$ and a random vector $X$, we say $X\in\mathcal{L}^{s}$ if $\lVert X\rVert_{s}=[\mathbb{E}(|X|^{s})]^{1/s}<\infty$. For two positive number sequences $(a_{n})$ and $(b_{n})$, we say $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ (resp. $a_{n}\asymp b_{n}$) if there exists $C>0$ such that $a_{n}/b_{n}\leq C$ (resp. $1/C\leq a_{n}/b_{n}\leq C$) for all large $n$, and say $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\rightarrow 0$ as $n\rightarrow\infty$. We set $(X_{n})$ and $(Y_{n})$ to be two sequences of random variables. Write $X_{n}=O_{\mathbb{P}}(Y_{n})$ if for any $\epsilon>0$, there exists $C>0$ such that $\mathbb{P}(X_{n}/Y_{n}\leq C)>1-\epsilon$ for all large $n$, and say $X_{n}=o_{\mathbb{P}}(Y_{n})$ if $X_{n}/Y_{n}\rightarrow 0$ in probability as $n\rightarrow\infty$.
@@ 2.1 Model Setup
In this subsection, we formulate our model. Recall the full model in (1.1) where $U_{jt}\in\mathbb{R}^{d}$ is a latent random vector that can be arbitrarily correlated with $X_{jt},$ $\epsilon_{jt}$ is a mean zero term that is independent of $\left(X_{jt},U_{jt}\right)$.
Without loss of generality set $c_{0j}=0$ for the known case. The full model in (1.1) can be rewritten into the following reduced form model:
$Y_{jt}=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt},$ (2.1)
where $\tilde{\tau}_{j}(X_{jt})=\mathbb{E}[\tau_{j}(X_{jt},U_{jt})|X_{jt}]$, $\tilde{h}_{j}(X_{jt})=\mathbb{E[}h_{j}(X_{jt},U_{jt})|X_{jt}]$, $\varepsilon_{j}(X_{jt},U_{jt})=h_{j}(X_{jt},U_{jt})-\tilde{h}_{j}(X_{jt})+[\tau_{j}(X_{jt},U_{jt})-\tilde{\tau}_{j}(X_{jt})]\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}$ and the reduced form error
@@
The noise term $e_{jt}$ has a complex structure, influenced by both the latent variable and by the heterogeneous volatility of the observations. This intricate structure, combined with the temporal and cross-sectional dependencies of the noise, makes the inference process particularly challenging.
A few remarks are in order. First, we have $\mathbb{E}(e_{jt}|X_{jt})=0$ and the conditional expectation of the observed outcome given $X_{jt}$ is $\mathbb{E}(Y_{jt}|X_{jt})=\tilde{h}_{j}(X_{jt})+\tilde{\tau}_{j}(X_{jt})\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}.$ Second, in Section G.2 we show the possibilities of extending our model to add more covariates and incoporating fixed effects. Third, ($X_{jt},U_{jt}$) can be lagged variables. For example, in our application on modeling stock volatility (Section 5.1), $X_{jt}=Y_{j(t-1)}$, functions $h_{j}(\cdot,\cdot)$, $\tau_{j}(\cdot,\cdot)$ and $\sigma_{j}(\cdot,\cdot)$ represent baseline effect, the jump effect and the volatility function, respectively. The $U_{jt}$ are unobserved risk factors or variables we exclude in the nonparametric regression. Modeling $U_{jt}$ is important in this circumstance as the volatility of a single stock is most likely affected by other risk factors such as news events or unobserved market factors.
This paper aims to explore methods for conducting simultaneous inference on threshold effect estimators in a large $N$ and large $T$ setup. In particular, we allow $N\gg T.$ Our focus is on the threshold effect, starting with the case where the threshold value $c_{0j}$ is known and then extending the analysis to the unknown case in Subsections 2.3 and 3.3.
Assuming that $X_{jt}$ is a continuous random variable, the jump effects can be identified as a “gap” in the conditional expectations and their derivatives at the thresholds. Let $\mu_{j}(x):=\mathbb{E}(Y_{jt}|X_{jt}=x)$ and $\mu_{j}(c_{0j}-)=\lim_{x\uparrow c_{0j}}\mu_{j}(x)$ and $\mu_{j}(c_{0j}+)=\lim_{x\downarrow c_{0j}}\mu_{j}(x).$ Consider $\partial_{+}\mu_{j}(\cdot)$ (resp. $\partial_{-}\mu_{j}(\cdot)$) as the right (resp. left) derivative of function $\mu_{j}(\cdot).$ The threshold effect for the individual $j$ is defined as
$\gamma_{j}$ $=$ $\mu_{j}(c_{0j}+)-\mu_{j}(c_{0j}-),$ (2.3)
$\gamma_{j}^{[1]}$ $=$ $\partial_{+}\mu_{j}(c_{0j})-\partial_{-}\mu_{j}(c_{0j}).$ (2.4)
Due to continuity of function $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$, we have $\gamma_{j}=\tilde{\tau}_{j}(c_{0j}),\gamma_{j}^{[1]}=\tilde{\tau}^{\prime}_{j}(c_{0j}).$ ($\tilde{\tau}^{\prime}_{j}(.)$ denotes the derivative of $\tilde{\tau}_{j}(.)$.)